### v1

@@
Abstract
@@
Index Terms:
Room acoustics, room impulse response estimation, sound field reconstruction, SIREN, residual connections, three-dimensional sound field.
[…]
=-15pt
I Introduction
@@
The application of residual connections to PINNs has also been explored to solve inverse problems. For instance, Tian et al. performed wind-field reconstruction using PINNs with an MLP augmented by residual connections [36]. This method achieved robust estimation under noisy conditions, and by mitigating the vanishing gradient problem, it enabled the accurate computation of the governing physical laws.
In addition, residual connections in SIREN-based architectures have been studied to facilitate deeper networks. Lu et al. demonstrated that adding residual connections to SIREN stabilized the training and improved the accuracy of the 3D image reconstruction [42]. Based on this architecture, we previously proposed relatively shallow PINNs incorporating SIREN with residual connections for 2D sound fields [43].
However, for inverse problems in PINN-based RIR estimation, the effect of network depth in architectures with residual connections has not been systematically studied. In addition, the relative suitability of activation functions (e.g., tanh vs. sin) for this acoustic-inverse problem remains unclear because direct comparisons are limited. Consequently, it is not yet well understood how increasing depth influences RIR estimation performance and which activation function is most appropriate for the task.
In this study, we investigated the acoustic-inverse problem of RIR estimation in a three-dimensional sound field to address these open questions. To this end, we employed a PINN architecture that combines SIREN with residual connections. We further conducted a direct comparison between SIREN-based and tanh-based PINN architectures across different depths, with and without residual connections, to quantify the effects of activation functions, network depth, and residual connections on RIR estimation performance.
@@ II Problem Formulation
III-A Physics-informed neural networks (Conventional Method)
@@ III-A2 Network Architecture
III-B PINNs with residual connections (Proposed Method)
@@
$f_{\mathrm{{PIRes}}}({\mathbf{w,x}})=(H_{{R}}\circ H_{R-1}\circ\cdots H_{{r}}\circ\cdots H_{1})(\mathbf{x}),$ (10)
@@ IV Experimental Setup
Figure 5: The trajectory of the loss function and estimation accuracy of evaluation and measured points. (a) The trajectory of the loss function $\mathcal{L}$ over iterations for models with $N=6$ hidden layers. (b) The trajectory of the loss function $\mathcal{L}$ over iterations for models with $N=18$ hidden layers. (c) NMSE of evaluation points versus the number of hidden layers. (d) NMSE of measurement points versus the number of hidden layers. (e) Frequency-domain NMSE for the best-performing model of each method, where the network depth is selected to maximize performance.
IV-A Simulation Setup
@@
RIRs were simulated using the image source method [45] with the SFS-toolbox [46], considering reflections up to the fourth order. The simulated signals had a length of 400 samples and a sampling frequency of 8000 Hz. To simulate realistic measurement conditions, white noise was added to the measured signals to achieve a signal-to-noise ratio (SNR) of 20 dB.
@@ IV-B Network parameters
IV-C evaluation method
@@
V Results
V-A Evaluation of overall estimation accuracy
@@
V-A1 Training Stability and Convergence
@@
Within each activation function, incorporating residual connections led to faster and more stable convergence. This suggests that residual connections can accelerate training and improve the convergence behavior even for shallow networks. In particular, comparing PINNs-SIREN with PINNs-SIREN-Res, the residual variant achieved a slightly lower final loss after $50,000$ iterations, indicating that the benefit of the residual connections is observable even at $N=6$. Overall, the proposed method achieved faster convergence and a lower final loss than its non-residual counterpart at shallow depth.
Fig. 5(b) shows the results for the deeper models ($N=18$). The total loss $\mathcal{L}$ converges for all methods. Consistent with the $N=6$ case, PINNs-SIREN-Res achieved the fastest and most stable convergence, reaching the lowest loss among the compared models. In addition, within each activation function, the residual variants converged faster and attained lower final losses than their non-residual counterparts. Overall, the convergence-speed trends observed for $N=6$ are preserved at $N=18$.
From these results, the loss converged for all the methods, even for a deeper setting ($N=18$). Across both depths, the SIREN-based models consistently converged faster and to lower loss values than the tanh-based models, and incorporating residual connections further improved the convergence, particularly in terms of speed and final loss.
TABLE I: NMSE [dB] of the lowest-NMSE model across the tested depths for each method with $50,000$ training iterations. This table presents the Normalized Mean Squared Error (NMSE) values achieved after $50,000$ iterations for each model (PINNs-tanh, PINNs-tanh-Res, PINNs-SIREN, and PINNs-SIREN-Res).
PINNs-tanh PINNs-tanh-Res PINNs-SIREN PINNs-SIREN-Res
2.1-2.1 7.8-7.8 14.6-14.6 16.3-16.3
TABLE II: NMSE [dB] in each frequency band for the evaluated models.
@@
010000-1000 8.2-8.2 13.2-13.2 23.2-23.2 24.6-24.6
100020001000-2000 1.5-1.5 6.7-6.7 15.5-15.5 17.9-17.9
200030002000-3000 0.6-0.6 6.3-6.3 12.8-12.8 14.5-14.5
300040003000-4000 0.3-0.3 5.7-5.7 10.1-10.1 11.5-11.5
V-A2 Estimation Accuracy over the Entire Evaluation Region
@@
To clarify the effect of network depth on estimation performance, we compare the mean NMSE across all evaluation points. Fig. 5(c) shows the relationship between NMSE and the number of hidden layers ($N=6,10,14,18$), and Table I lists the NMSE of the lowest-NMSE model across the tested depths for each method.
@@
First, for the tanh-based models, PINNs-tanh showed little improvement as depth increased: the NMSE was approximately 1.0-1.0 dB at $N=6$ and 2.7-2.7 dB at $N=18$. In contrast, PINNs-tanh-Res improved markedly with depth, reaching an overall NMSE of approximately 7.9-7.9 dB at $N=18$, which is a 5.2 dB improvement over PINNs-tanh. Moreover, increasing the depth from $N=6$ to $N=18$ improved the NMSE of PINNs-tanh-Res by approximately $6.0$ dB. These results indicate that residual connections are important for effectively exploiting increased depth in tanh-based PINNs.
@@
Second, the results indicate that PINNs-SIREN achieves higher estimation accuracy than the tanh-based models among the compared models. For the best-performing models (across depths) of each type, PINNs-SIREN improved the NMSE by approximately $12.5$ dB compared with PINNs-tanh and by $6.8$ dB compared with PINNs-tanh-Res. As shown in Fig. 5(c), the NMSE of PINNs-SIREN improved as the depth increased up to $N=14$; however, further increasing the depth to $N=18$ led to a slight decrease in performance (approximately $0.6$ dB). This suggests that PINNs-SIREN outperforms tanh-based models; however, its estimation accuracy may degrade beyond a certain depth.
@@
Next, we evaluate the proposed method, PINNs-SIREN-Res. As summarized in Table I, the best-performing model (across depths) achieves NMSE improvements of approximately $14.2$ dB over PINNs-tanh, $8.5$ dB over PINNs-tanh-Res, and $1.7$ dB over PINNs-SIREN. Notably, while PINNs-SIREN showed a slight degradation when the depth was increased from $N=14$ to $N=18$, the proposed method instead improved the NMSE by 0.7 dB. These results indicate that the best-performing proposed model outperforms the corresponding best-performing conventional models, and that its estimation accuracy improves with increasing depth up to $N=18$.
@@
To assess the noise-reduction effect, Fig. 5(d) shows the mean NMSE at the measurement points. For all methods except PINNs-SIREN, the estimation accuracy improved with increasing depth. In contrast, PINNs-SIREN shows a decrease in accuracy when the network is deepened to $N=18$. The proposed method improved the NMSE by approximately $17.6$ dB compared with PINNs-tanh, $12.0$ dB compared with PINNs-tanh-Res, and $2.9$ dB compared with PINNs-SIREN. Consistent with the trends observed at the evaluation points, these results indicate that the proposed method achieved the highest estimation accuracy at the measurement positions.
@@
TABLE III: NMSE [dB] in the interpolation and extrapolation regions (inside and outside the measurement sphere, respectively) for each model.
@@
Interpolation 2.5-2.5 7.9-7.9 15.6-15.6 17.8-17.8
Extrapolation 2.3-2.3 7.8-7.8 14.0-14.0 15.5-15.5
V-A3 Frequency-Domain Accuracy
@@
Table II summarizes the NMSE in four frequency bands: up to 1000 Hz, 1000–2000 Hz, 2000–3000 Hz, and 3000–4000 Hz.
PINNs-tanh attains an NMSE of -8.2 dB for frequencies up to 1000 Hz; however, its performance degrades substantially in higher bands compared with the other methods. For PINNs-tanh-Res, the NMSE increases by approximately 6.5 dB from the sub-1000 Hz band to the 1000–2000 Hz band, but remains around 6-6 dB above 2000 Hz.
PINNs-SIREN achieves markedly better accuracy up to 1000 Hz (NMSE: 23.2-23.2 dB), but its NMSE gradually increases above 1000 Hz. The proposed PINNs-SIREN-Res achieves the best NMSE up to 1000 Hz (24.6-24.6 dB) and shows less degradation than PINNs-SIREN at higher frequencies, indicating more stable estimation in the high-frequency bands.
@@
V-B Evaluation at Specific points
@@
V-B1 Interpolation and Extrapolation Performance
@@
Fig. 6(a) corresponds to the plane at z=0.0115=-0.0115 m, which is the central slice of the evaluation domain along the z-axis. Note that, on this x–y plane, the measurement sphere is close to the boundary of the evaluation domain; consequently, many measurement points are concentrated near the periphery of the evaluation domain. Therefore, this slice is representative of interpolation performance. As shown in the figure, PINNs-tanh and PINNs-tanh-Res generally yielded a higher NMSE (i.e., lower accuracy) than the SIREN-based models across the evaluation points. The proposed method reduces the NMSE by approximately $14.9$ dB, $9.9$ dB, and $2.2$ dB compared with PINNs-tanh, PINNs-tanh-Res, and PINNs-SIREN, respectively.
@@
TABLE IV: NMSE [dB] for the early and late segments of the RIR. This table lists the NMSE for each model for the early segment (direct sound and early reflections) and the late segment (late reverberation). The late segment is defined as the portion of the RIR after $0.025$ s.
Early segment 2.4-2.4 9.3-9.3 18.0-18.0 19.0-19.0
Late segment 1.8-1.8 2.3-2.3 6.9-6.9 9.2-9.2
@@
Table III summarizes the NMSE for the evaluation points located inside (interpolation region) and those outside (extrapolation region) the measurement points. In the interpolation region, the proposed method reduced the NMSE by approximately $15.3$ dB relative to PINNs-tanh, $9.9$ dB relative to PINNs-tanh-Res, and $2.2$ dB relative to PINNs-SIREN. In the extrapolation region, it reduced the NMSE by approximately $13.2$ dB, $7.7$ dB, and $1.5$ dB compared with these methods, respectively. These results suggest that incorporating residual connections improved the estimation accuracy in both interpolation and extrapolation.
V-B2 Time-Domain Analysis of Estimation Accuracy

### v2

@@
[…]
Spatial Interpolation of Room Impulse Responses based on Deeper Physics-Informed Neural Networks with Residual Connections
keywords:
Room acoustics, room impulse response estimation, sound field reconstruction, SIREN, residual connections
titlenote: This work was partially supported by JSPS KAKENHI Grant Number 22K12099 and JGC-Saneyoshi Scholarship Foundation. This paper has been submitted to a preprint server [47].
{summary}
@@
1 Introduction
@@
The application of residual connections to PINNs has also been explored to solve inverse problems. For instance, Tian et al. performed wind-field reconstruction using PINNs with an MLP augmented by residual connections [36]. This method achieved robust estimation under noisy conditions, and by mitigating the vanishing gradient problem, it enabled the accurate computation of the governing physical laws. However, it should be noted that the model employed in their study was not a SIREN-based MLP, but rather a standard MLP utilizing a tanh activation function.
In addition, residual connections in SIREN-based architectures have been studied to facilitate deeper networks. Lu et al. demonstrated that adding residual connections to SIREN without physical laws stabilized the training and improved the accuracy of the 3D image reconstruction [42]. Based on this architecture, we previously proposed relatively shallow PINNs incorporating SIREN with residual connections for 2D sound fields [43]. This previous model employed a compact architecture consisting of only 4 hidden layers with 256 neurons per layer. Consequently, the impact of hidden layer depth on estimation accuracy was not verified in that study, nor was estimation performed in a three-dimensional sound field.
Therefore, for inverse problems in PINN-based RIR estimation, the effect of network depth in architectures with residual connections has not been systematically studied. In addition, the relative suitability of activation functions (e.g., tanh vs. sin) for this acoustic-inverse problem remains unclear because direct comparisons are limited. Consequently, it is not yet well understood how increasing depth influences RIR estimation performance and which activation function is most appropriate for the task.
In this study, we investigate the acoustic-inverse problem of RIR estimation in a three-dimensional sound field to address these open questions. To this end, we consturuct a deeper PINN architecture that combines SIREN with residual connections. Furthermore, we quantitatively compare the effects of activation functions, network depth, and residual connections, providing insight into the design of deep and stable PINN for acoustic inverse problems.
The main contributions of this paper are summarized as follows:
  • Systematic depth study for 3D RIR PINNs: We conduct a controlled investigation of network depth in PINNs for 3D RIR estimation, clarifying when deeper models degrade or improve performance under physics-based training.
  • Residual SIREN tailored for physics-informed training: We propose a deep physics-informed architecture that combines SIREN activations with residual connections (PINNs-SIREN-Res) to stabilize training in deep settings and to improve reconstruction accuracy for RIR estimation.
  • Task-oriented evaluation and insights: Through extensive comparisons (tanh vs. sin, with/without residuals), we provide practical design insights validated on interpolation and extrapolation, temporal behavior, and frequency-band performance.
@@ 2 Problem Formulation
3.1 Physics-informed neural networks (Conventional Method)
@@ 3.1.2 Network Architecture
3.2 PINNs with residual connections (Proposed Method)
@@
$f_{\mathrm{{PIRes}}}({\mathbf{w,x}})=(H_{{R}}\circ H_{R-1}\circ\cdots H_{{r}}\circ\cdots H_{1})(\mathbf{x}),\\$ (10)
@@ 4 Experimental Setup
Figure 5: The trajectory of the loss function and estimation accuracy of evaluation and measured points. The trajectory of the loss function $\mathcal{L}$ over iterations for models with (a) $N=6$ and (b) $N=18$ hidden layers. (c) NMSE of evaluation points versus the number of hidden layers. (d) NMSE of measurement points versus the number of hidden layers. (e) Frequency-domain NMSE for the best-performing model of each method, where the network depth is selected to maximize performance.
4.1 Simulation Setup
@@
The simulated signals had a length of 400 samples and a sampling frequency of 8000 Hz. To simulate realistic measurement conditions, white noise was added to the measured signals to achieve a signal-to-noise ratio (SNR) of 20 dB. RIRs were simulated using the image source method [45] with the SFS-toolbox [46], considering reflections up to the seventh order based on these room conditions and signal length.
@@ 4.2 Network parameters
The training of all methods was performed on a workstation equipped with an AMD EPYC 7543P CPU (32 cores, 64 threads, 2.8 GHz), an NVIDIA RTX A6000 GPU (48 GB VRAM), and 256 GB of RAM.
4.3 evaluation method
@@
5 Results
Table 1: NMSE [dB] of the lowest-NMSE model across the tested depths for each method with $50,000$ training iterations. This table presents the Normalized Mean Squared Error (NMSE) values achieved after $50,000$ iterations for each model (PINNs-tanh, PINNs-tanh-Res, PINNs-SIREN, and PINNs-SIREN-Res).
PINNs-tanh PINNs-tanh-Res PINNs-SIREN PINNs-SIREN-Res
2.6-2.6 6.7-6.7 12.5-12.5 14.2-14.2
5.1 Evaluation of overall estimation accuracy
@@
5.1.1 Training Stability and Convergence
@@
Table 2: NMSE [dB] in each frequency band for the evaluated models.
@@
010000-1000 7.2-7.2 13.2-13.2 21.2-21.2 23.7-23.7
100020001000-2000 1.3-1.3 5.5-5.5 12.7-12.7 14.6-14.6
200030002000-3000 1.0-1.0 4.8-4.8 10.2-10.2 12.0-12.0
300040003000-4000 0.8-0.8 3.9-3.9 8.1-8.1 9.4-9.4
Within each activation function, incorporating residual connections led to faster convergence. This suggests that residual connections can accelerate training and improve convergence behavior even for shallow networks. In particular, comparing PINNs-SIREN with PINNs-SIREN-Res, the residual variant achieved a comparable final loss after $50,000$ iterations. This indicates that while residual connections accelerate training, they yield equivalent accuracy at $N=6$. Overall, the proposed method achieved faster convergence while maintaining a comparable final loss relative to its non-residual counterpart at shallow networks.
Fig. 5(b) shows the results for the deeper models ($N=18$). The total loss $\mathcal{L}$ successfully converged for all methods. Consistent with the $N=6$ case, PINNs-SIREN-Res achieved the fastest and most stable convergence, reaching the lowest loss among the compared models. Furthermore, within each activation function, the residual variants converged faster and attained lower final losses than their non-residual counterparts. Overall, the convergence trends observed at $N=6$ were preserved at $N=18$.
From these results, convergence was achieved for all methods, even in the deeper setting. Across both depths, the SIREN-based models consistently demonstrated faster convergence. Specifically, in the deeper setting, SIREN-based models achieved lower loss values than tanh-based models, and incorporating residual connections further improved performance, particularly in terms of convergence speed and the final loss.
5.1.2 Estimation Accuracy over the Entire Evaluation Region
@@
To clarify the effect of network depth on estimation performance, we compare the mean NMSE across all evaluation points. Fig. 5(c) shows the relationship between NMSE and the number of hidden layers ($N=6,10,14,18$), and Table 1 lists the NMSE of the lowest-NMSE model across the tested depths for each method.
@@
First, for the tanh-based models, PINNs-tanh showed little improvement as depth increased: the NMSE was approximately 1.1-1.1 dB at $N=6$ and 2.6-2.6 dB at $N=18$. In contrast, PINNs-tanh-Res improved markedly with depth, reaching an overall NMSE of approximately 6.7-6.7 dB at $N=18$, which is a 4.1 dB improvement over PINNs-tanh. Moreover, increasing the depth from $N=6$ to $N=18$ improved the NMSE of PINNs-tanh-Res by approximately $3.9$ dB. These results indicate that residual connections are important for effectively exploiting increased depth in tanh-based PINNs.
@@
Second, the results indicate that PINNs-SIREN achieves higher estimation accuracy than the tanh-based models among the compared models. For the best-performing models (across depths) of each type, PINNs-SIREN improved the NMSE by approximately $9.9$ dB compared with PINNs-tanh and by $5.8$ dB compared with PINNs-tanh-Res. As shown in Fig. 5(c), the NMSE of PINNs-SIREN improved as the depth increased up to $N=14$; however, further increasing the depth to $N=18$ led to a slight decrease in performance (approximately $0.2$ dB). This suggests that PINNs-SIREN outperforms tanh-based models; however, its estimation accuracy may degrade beyond a certain depth.
@@
Next, we evaluate the proposed method, PINNs-SIREN-Res. As summarized in Table 1, the best-performing model (across depths) achieves NMSE improvements of approximately $11.6$ dB over PINNs-tanh, $7.5$ dB over PINNs-tanh-Res, and $1.7$ dB over PINNs-SIREN. Notably, while PINNs-SIREN showed a slight degradation when the depth was increased from $N=14$ to $N=18$, the proposed method instead improved the NMSE by 0.4 dB. These results indicate that the best-performing proposed model outperforms the corresponding best-performing conventional models, and that its estimation accuracy improves with increasing depth up to $N=18$.
@@
To assess the noise-reduction effect, Fig. 5(d) shows the mean NMSE at the measurement points. For all methods except PINNs-SIREN, the estimation accuracy improved with increasing depth. In contrast, PINNs-SIREN shows a decrease in accuracy when the network is deepened to $N=18$. The proposed method improved the NMSE by approximately $14.9$ dB compared with PINNs-tanh, $10.5$ dB compared with PINNs-tanh-Res, and $3.0$ dB compared with PINNs-SIREN. Consistent with the trends observed at the evaluation points, these results indicate that the proposed method achieved the highest estimation accuracy at the measurement positions.
@@
Table 3: NMSE [dB] in the interpolation and extrapolation regions (inside and outside the measurement sphere, respectively) for each model.
@@
Interpolation 2.6-2.6 6.7-6.7 13.1-13.1 15.0-15.0
Extrapolation 2.5-2.5 6.7-6.7 12.1-12.1 13.7-13.7
5.1.3 Frequency-Domain Accuracy
@@
Table 2 summarizes the NMSE in four frequency bands: up to 1000 Hz, 1000–2000 Hz, 2000–3000 Hz, and 3000–4000 Hz.
PINNs-tanh attains an NMSE of -8.2 dB for frequencies up to 1000 Hz; however, its performance degrades substantially in higher bands compared with the other methods. For PINNs-tanh-Res, the NMSE increases by approximately $7.7$ dB from the sub-1000 Hz band to the 1000–2000 Hz band, but remains around 5-5 dB above 2000 Hz.
PINNs-SIREN achieves markedly better accuracy up to 1000 Hz (NMSE: 21.2-21.2 dB), but its NMSE gradually increases above 1000 Hz. The proposed PINNs-SIREN-Res achieves the best NMSE up to 1000 Hz (23.7-23.7 dB) and shows less degradation than PINNs-SIREN at higher frequencies, indicating more stable estimation in the high-frequency bands.
@@
Table 4: NMSE [dB] for the early and late segments of the RIR. This table lists the NMSE for each model for the early segment (direct sound and early reflections) and the late segment (late reverberation). The late segment is defined as the portion of the RIR after $0.025$ s.
Early segment 2.5-2.5 8.3-8.3 17.1-17.1 17.5-17.5
Late segment 2.8-2.8 3.5-3.5 7.3-7.3 9.6-9.6
5.2 Evaluation at Specific points
@@
5.2.1 Interpolation and Extrapolation Performance
@@
Fig. 6(a) corresponds to the plane at z=0.0115=-0.0115 m, which is the central slice of the evaluation domain along the z-axis. Note that, on this x–y plane, the measurement sphere is close to the boundary of the evaluation domain; consequently, many measurement points are concentrated near the periphery of the evaluation domain. Therefore, this slice is representative of interpolation performance. As shown in the figure, PINNs-tanh and PINNs-tanh-Res generally yielded a higher NMSE (i.e., lower accuracy) than the SIREN-based models across the evaluation points. The proposed method reduces the NMSE by approximately $11.5$ dB, $8.1$ dB, and $2.0$ dB compared with PINNs-tanh, PINNs-tanh-Res, and PINNs-SIREN, respectively.
@@
Table 3 summarizes the NMSE for the evaluation points located inside (interpolation region) and those outside (extrapolation region) the measurement points. In the interpolation region, the proposed method reduced the NMSE by approximately $12.4$ dB relative to PINNs-tanh, $8.3$ dB relative to PINNs-tanh-Res, and $1.9$ dB relative to PINNs-SIREN. In the extrapolation region, it reduced the NMSE by approximately $11.2$ dB, $7.0$ dB, and $1.6$ dB compared with these methods, respectively. These results suggest that incorporating residual connections improved the estimation accuracy in both interpolation and extrapolation.
5.2.2 Time-Domain Analysis of Estimation Accuracy